# Limit Computation Over Posets via Minimal Initial Functors

### v1

@@ 9. Questions for Future Work
We imagine that it may be possible to fully describe and efficiently compute minimal initial functors valued in arbitrary very small categories, extending our results about posets. It would also be worthwhile to study the optimality of our algorithmic results, particularly for limit computation. Finally, given that limits and diagrams of vectors spaces indexed by posets are both ubiquitous in mathematics, we are hopeful that the results of this paper could find practical use in applications beyond the generalized rank computations considered here. The development of such applications remains a subject for future work.

### v2

@@ 9. Questions for Future Work
We imagine that it may be possible to fully describe and efficiently compute minimal initial functors valued in arbitrary very small categories, extending our results about posets. It would also be worthwhile to study the optimality of our algorithmic results, particularly for limit computation. The algorithms of this paper for computing limits and generalized ranks should be relatively easy to implement using ordinary (cubic-time) Gaussian elimination; a good implementation would open the door to applications. In particular, it would be interesting to explore applications to the featurization of 3-parameter persistent homology for supervised learning, along the lines of what Xin et al. [41] have done in the the 2-parameter case.